# Unifying Decoherence and Phase Evolution in Mixed Quantum–Classical Dynamics through Exact Factorization

### v1

@@ Abstract
Figure 3: Spatial distribution of nuclear wave packets ($|\chi_{i}(R,t)|^{2}$) and BO populations ($|C_{i}(R,t)|^{2}$) at different time steps, $t=1300$ a.u. (a,c,e,g) and $1900$ a.u. (b,d,f,h) with the initial momentum $k_{0}=50.0$ a.u. for CT-based approaches (CT/CTv2) (a-d) and SHXF-based approaches (SHXF/SHXFv2) (e-h). The BO potential energies $\epsilon_{1/2}$ and the NAC $d_{12}$ are depicted to interpret wave-packet dynamics (a,b,e,f).
@@
In addition, to analyze the wave-packet dynamics, we construct the spatial distribution of the BO-projected time-dependent nuclear densities, $|\chi_{i}(R,t)|^{2}$, and BO populations, $|C_{i}(R,t)|^{2}$, for the DAG model. $|\chi_{i}(R,t)|^{2}$ are obtained from histograms of nuclear trajectories weighted by their corresponding BO populations, while $|C_{i}(R,t)|^{2}$ are calculated by $|C_{i}(R,t)|^{2}=|\chi_{i}(R,t)|^{2}/|\chi(R,t)|^{2}$ where the total nuclear density, $|\chi(R)|^{2}$, is the histograms of trajectories. Fig. 3 presents $|\chi_{i}(R,t)|^{2}$ and $|C_{i}(R,t)|^{2}$ from the XF-based approaches for large $k_{0}$ ($k_{0}=50$ a.u.), at $t=1300$ a.u. (arrival of nuclear wave packets at the second NAC region) and $t=1900$ a.u. (after passage through the second NAC region). We observe a marked improvement in the spatial distributions with CTv2 and SHXFv2. CT exhibits oscillatory behavior arising from incorrect phase estimation in each BO state, while SHXF yields an overall inaccurate spatial distribution. By contrast, CTv2 reproduces almost identical results to the exact dynamics. SHXFv2 performs reasonably well at $k=1900$ a.u., but at $k=1300$ a.u. it produces strongly overlapping nuclear wave packets, likely due to a less accurate treatment of decoherence.
@@
Figure 4: Spatial distribution of nuclear wave packets (a,b) and BO populations (c,d) at $t=2200$ a.u. with the initial momentum $k_{0}=30.0$ a.u. for CTv2 (a,c) and CTv2 with the QM $\mathbf{\mathcal{P}}_{\nu}$ and the PQM $\mathbf{\mathcal{Q}}_{\nu}$ obtained from quantum dynamics (b,d). BO potentials and NAC are depicted as in Fig. 3.
@@
For $k_{0}=30.0$ a.u., the nuclear wave-packet splitting and the electronic decoherence effects are pronounced after passing the first NAC region because of the small kinetic energy, due to a larger relative difference in the velocities of nuclear wave packets on different BO states. Thus, the well-separated nuclear wave packets pass the second NAC region sequentially. Fig. 4 shows nuclear wave packets and BO populations at $t=2200$ a.u. when $|\chi_{1}(R,t)|^{2}$ meets the second NAC region. CTv2 provides better wave-packet splitting and $|C_{i}(R,t)|^{2}$ compared to CT in Fig. LABEL:fig:ct_spatial_k0_30 [33], while both deviate from the exact result. On the other hand, the CTv2 with the exact QM and PQM recovers the correct $|\chi_{i}(R,t)|^{2}$ and $|C_{i}(R,t)|^{2}$, whereas CT [Fig. LABEL:fig:ct_spatial_k0_30] with the exact QM still deviates from the exact result [33]. Thus, the formulation is valid; the discrepancies mainly originate from the approximations to $\bm{\mathcal{P}}_{\nu}$ and $\bm{\mathcal{Q}}_{\nu,ik}$, as the true $|\chi_{2}(R,t)|^{2}$ becomes strongly non-Gaussian after the first NAC region for $k_{0}=30.0$ a.u.
@@
The results from other approaches are provided in Figs. LABEL:fig:ct_spatial_k0_30LABEL:fig:pcsh_both_spatial_k0_50 [33]. For all $k_{0}$, PCSH and PCSH-EDC yields either incorrect spatial decoherence or incorrect BO population after the second NAC region, respectively, whereas CTv2 and SHXFv2 exhibit clear enhancement compared to the other methods.
Figure 5: Transmission probability for the lower state ($T_{1}$) for the 2DNS model, as in Fig. 1.
@@
Finally, to test the additional corrections in higher-dimensional systems, we compute the $T_{1}$ for the 2DNS model. As shown in Fig. 5, SHXFv2 and CTv2, along with PCSH, reproduce the correct oscillation pattern, whereas PCSH-EDC and SHXF fail, consistent with the DAG results. CT works for $k_{0}>25$ a.u. but fails for $k_{0}=17$–25 a.u. Thus, the phase and PQM corrections enable us to reproduce the quantum dynamics even in higher dimensions.

### v2

@@ Abstract
We also present the transmission probability $T_{1}$ and the time evolution of BO populations determined by the active states of the trajectories to test the internal consistency of SHXFv2, in Figures LABEL:fig:RT_run and LABEL:fig:pop_run of the Supporting Information, respectively. These results are in good agreement with those calculated from the electronic states of the trajectories (Figures 1 and 2), suggesting that both SHXF and SHXFv2 exhibit internal consistency. However, as previously noted, SHXF fails to accurately capture the transmission probability and BO population after passing through the second NAC region despite its internal consistency.
@@
Figure 3: Spatial distribution of nuclear wave packets ($|\chi_{i}(R,t)|^{2}$) (a,b,e,f) and BO populations ($|C_{i}(R,t)|^{2}$) (c,d,g,h) at different time steps, $t=1300$ a.u. (a,c,e,g) and $1900$ a.u. (b,d,f,h) with the initial momentum $k_{0}=50.0$ a.u. for CT-based approaches (CT/CTv2) (a-d) and SHXF-based approaches (SHXF/SHXFv2) (e-h). The BO potential energies $\epsilon_{1/2}$ and the NAC $d_{12}$ are depicted to interpret wave-packet dynamics (a,b,e,f).
@@
In addition, to analyze the wave-packet dynamics, we construct the spatial distribution of the BO-projected time-dependent nuclear densities, $|\chi_{i}(R,t)|^{2}$, and BO populations, $|C_{i}(R,t)|^{2}$, for the DAG model. $|\chi_{i}(R,t)|^{2}$ are obtained from histograms of nuclear trajectories weighted by their corresponding BO populations, while $|C_{i}(R,t)|^{2}$ are calculated by $|C_{i}(R,t)|^{2}=|\chi_{i}(R,t)|^{2}/|\chi(R,t)|^{2}$ where the total nuclear density, $|\chi(R)|^{2}$, is the histograms of trajectories. Figure 3 presents $|\chi_{i}(R,t)|^{2}$ and $|C_{i}(R,t)|^{2}$ from the XF-based approaches for large $k_{0}$ ($k_{0}=50.0$ a.u.), at $t=1300$ a.u. (arrival of nuclear wave packets at the second NAC region) and $t=1900$ a.u. (after passage through the second NAC region). We observe a marked improvement in the spatial distributions with CTv2 and SHXFv2. CT exhibits oscillatory behavior arising from incorrect phase estimation in each BO state, while SHXF yields an overall inaccurate spatial distribution. By contrast, CTv2 reproduces almost identical results to the exact dynamics. SHXFv2 performs reasonably well at $t=1900$ a.u., but at $t=1300$ a.u. it produces strongly overlapping nuclear wave packets, likely due to a less accurate treatment of decoherence.
@@
Figure 4: Spatial distribution of nuclear wave packets (a,b) and BO populations (c,d) at $t=2200$ a.u. with the initial momentum $k_{0}=30.0$ a.u. for CTv2 (a,c) and CTv2 with the QM $\mathbf{\mathcal{P}}_{\nu}$ and the PQM $\mathbf{\mathcal{Q}}_{\nu,kl}$ obtained from quantum dynamics (b,d). BO potentials and NAC are depicted as in Figure 3.
@@
For $k_{0}=30.0$ a.u., the nuclear wave-packet splitting and the electronic decoherence effects are pronounced after passing the first NAC region due to a larger relative difference in the velocities of nuclear wave packets on different BO states, because of the small initial kinetic energy. Thus, the well-separated nuclear wave packets pass the second NAC region sequentially. Figure 4 shows nuclear wave packets and BO populations at $t=2200$ a.u. when $|\chi_{1}(R,t)|^{2}$ meets the second NAC region. CTv2 provides better wave-packet splitting and $|C_{i}(R,t)|^{2}$ compared to CT in Figure LABEL:fig:ct_spatial_k0_30, while both deviate from the exact result. On the other hand, the CTv2 with the exact QM and PQM recovers the correct $|\chi_{i}(R,t)|^{2}$ and $|C_{i}(R,t)|^{2}$, whereas CT (Figure LABEL:fig:ct_spatial_k0_30) with the exact QM still deviates from the exact result. Thus, the formulation is valid; the discrepancies mainly originate from the approximations to $\bm{\mathcal{P}}_{\nu}$ and $\bm{\mathcal{Q}}_{\nu,ik}$, as the true $|\chi_{2}(R,t)|^{2}$ becomes strongly non-Gaussian after the first NAC region for $k_{0}=30.0$ a.u.
@@
The results from other approaches are provided in Figures LABEL:fig:ct_spatial_k0_30LABEL:fig:pcsh_both_spatial_k0_50. For all $k_{0}$, PCSH and PCSH-EDC yields either incorrect spatial decoherence or incorrect BO population after the second NAC region, respectively, whereas CTv2 and SHXFv2 exhibit clear enhancement compared to the other methods.
Regarding the relative importance of the ENC terms, the electron-nuclear correlation effects can be separated into two groups governing the amplitude and phase evolutions of the BO coefficient, namely the decoherence ($\dot{C}^{\mathrm{QM}}_{j}$, $\dot{C}^{\mathrm{PQM}}_{j}$, and $\dot{C}^{\mathrm{Div}}_{j}$) and phase correction ($\dot{C}^{\mathrm{Ph}}_{j}$) terms, respectively. Also, considering the order of terms in $\hbar$, the phase correction would have much larger magnitude than the decoherence correction terms. Therefore, we analyze the decoherence and phase correction terms separately. For the phase correction term, we analyze the relative phase evolution. We show the time evolution of each component of $\langle\dot{\rho}_{11}\rangle$ and $\langle\Delta\dot{S}_{12}\rangle$ (Figure 5), and the spatial characters of $F$, $\dot{\rho}_{11}$, and $\dot{S}_{12}$ at tt=950 a.u. and 1250 a.u. (Figure LABEL:fig:element_spatial) for $k_{0}=40.0$, where $\dot{\rho}^{\mathrm{QM/PQM/Div}}_{11}=2\Re(\dot{C}^{\mathrm{QM/PQM/Div}}_{1}C^{*}_{1})$ represents the population change induced by each ENC component and $\Delta\dot{S}^{\mathrm{Ph/BO}}_{12}=\dot{S}^{\mathrm{Ph/BO}}_{1}-\dot{S}^{\mathrm{Ph/BO}}_{2}$ is the relative phase change from $C^{\mathrm{Ph}}_{j}$ and the BO energy contribution in $C^{\mathrm{Eh}}_{j}$.
Figure 5: Time evolution of the individual components of the averaged time derivative of BO population for state 1 $\langle\dot{\rho}^{\mathrm{QM/PQM/Div}}_{11}\rangle(t)$ (upper) and the phase difference $\langle\Delta\dot{S}^{\mathrm{Ph/BO}}_{12}\rangle(t)$ (lower). Results obtained from CTv2 are shown in comparison with exact quantum wave packet dynamics for $k_{0}=40.0$ a.u.
Overall, the CTv2 results well reproduce the exact results while the spatial shape of the divergence term contribution is averaged out across the trajectories due to the approximation in the numerical implementation. The PQM and divergence contributions have magnitudes comparable to those of the QM term, while their temporal evolution and spatial characteristics differ from the QM contribution. Meanwhile, the phase correction effect is much smaller in magnitude than the phase evolution effect from the BO energy: $\Delta\dot{S}^{\mathrm{BO}}_{12}=-(\epsilon_{1}-\epsilon_{2})$. However, its effect is non-negligible as seen in Figure 1, because the phase evolution is more sensitive than the magnitude of the BO coefficient. Therefore, these newly introduced terms are essential for capturing the correct electron-nuclear correlation effect on both decoherence and phase evolutions.
In addition, we investigate the independent implementation of the PQM or phase term in SHXF and CT in Figures LABEL:fig:RT_pqm_phLABEL:fig:pop_coh_pqm_phLABEL:fig:ct_pqm_ph_spatial_k0_40, and LABEL:fig:shxf_pqm_ph_spatial_k0_40. The results demonstrate that incorporating only the PQM term improves the description of intermediate decoherence but fails to capture the correct population dynamics when passing through the second NAC region. Conversely, including only the phase term yields accurate population dynamics at the second NAC region but cannot fully describe the intermediate decoherence. Thus, the full consideration is essential to provide correct dynamics.
Figure 6: Transmission probability for the lower state ($T_{1}$) for the 2DNS model, as in Figure 1.
@@
Finally, to test the additional corrections in higher-dimensional systems, we compute the $T_{1}$ for the 2DNS model. As shown in Figure 6, SHXFv2 and CTv2, along with PCSH, reproduce the correct oscillation pattern, whereas PCSH-EDC and SHXF fail, consistent with the DAG results. CT works for $k_{0}>25$ a.u. but fails for $k_{0}=17$–25 a.u. Thus, the phase and PQM corrections enable us to reproduce the quantum dynamics even in higher dimensions.
Regarding the numerical stability of CTv2 and SHXFv2, we investigate the mean absolute errors (MAEs) for norm conservation of the SHXFv2 and CTv2 simulations for $k_{0}=40.0$ a.u. for the DAG model. (Figure LABEL:fig:mae) Both SHXFv2 and CTv2 exhibit reliable numerical stability comparable to that of SHXF and CT methods (MAE $10^{-8}$). Although CTv2 and SHXFv2 involve additional terms from $\hat{H}^{(2)}_{ENC}$, each term from eqs 5 and 9 can be obtained from the same ingredients required for CT and SHXF without any additional quantum chemical calculations. Consequently, the computational complexity of these new versions remains the same as the original methods.